# NexusFlow: Unifying Disparate Tasks under Partial Supervision via Invertible Flow Networks

### v1

@@ 4.2 N-Task Challenge on NYU-V2
Dataset and PS-MTL Protocol. Although our main experiments focus on the two-task setting, we further evaluate NexusFlow under a more general $N$-task MTL framework. Due to constraints on data availability, we consider a simpler homogeneous dense-prediction setup and adopt the NYU-V2 dataset [silberman2012indoor], which provides three standard tasks: 13-class semantic segmentation, depth estimation, and surface normal prediction. Following common practice, all images are resized to $288\times 384$ [liu2019end]. To simulate the PS-MTL setting, instead of randomly selecting a single task label per sample and masking the losses of the others [ruder2019latent, li2022learning], we partition the entire dataset into three roughly equal subsets using clustering, and assign exactly one task’s ground-truth label to each subset. This enforces a realistic partial supervision regime in which only one task annotation is available per training example.
Baselines and Architecture. We adopt a standard MTL architecture, which is common for homogeneous dense-prediction tasks [li2022learning]. This consists of a Multi-Task Attention Network [liu2019end], including a feature encoder followed by three lightweight, task-specific decoders. We compare NexusFlow against several key baselines: (1) Oracle, trained with full supervision for all three tasks. (2) Baseline, representing the standard shared-encoder model trained only on our PS-MTL Protocol. (3) MTPSL [li2022learning] with our PS-MTL protocol.
@@
Implementation Details. For the NYU-V2 experiments, all methods are trained from scratch following the default setting in MTPSL [li2022learning] on one NVIDIA A100 80G GPU. Our proposed NexusFlow module, implemented with $N=3$ invertible coupling layers, is inserted between the feature encoder and the three task decoders to align their latent feature distributions.
Quantitative Results. We present our results in Table 4. Our proposed NexusFlow significantly outperforms all partially supervised baselines. Although still left behind with the fully supervised Oracle, it has a strong advantage in computational cost, especially when compared with MTPSL [li2022learning] in terms of training time and GPU memory (shown in Table 5).

### v2

@@ 4.2 N-Task Challenge on NYU-V2
Dataset and PS-MTL Protocol. Although our main experiments focus on the two-task setting, we further evaluate NexusFlow under a more general $N$-task MTL framework. Due to constraints on data availability, we consider a simpler homogeneous dense-prediction setup and adopt the NYU-V2 dataset [40], which provides three standard tasks: 13-class semantic segmentation, depth estimation, and surface normal prediction. Following common practice, all images are resized to $288\times 384$ [34]. To simulate the PS-MTL setting, instead of randomly selecting a single task label per sample and masking the losses of the others [39, 31], we partition the entire dataset into three roughly equal subsets using clustering, and assign exactly one task’s ground-truth label to each subset. This enforces a realistic partial supervision regime in which only one task annotation is available per training example.
Baselines and Architecture. We adopt a standard MTL architecture, which is common for homogeneous dense-prediction tasks [31]. This consists of a Multi-Task Attention Network [34], including a feature encoder followed by three lightweight, task-specific decoders. We compare NexusFlow against several key baselines: (1) Full-supervision , trained with full supervision for all three tasks. (2) Baseline, representing the standard shared-encoder model trained only on our PS-MTL Protocol. (3) MTPSL [31] with our PS-MTL protocol.
@@
Implementation Details. For the NYU-V2 experiments, all methods are trained from scratch following the default setting in MTPSL [31] on one NVIDIA A100 80G GPU. Our proposed NexusFlow module, implemented with $N=3$ invertible coupling layers, is inserted between the feature encoder and the three task decoders to align their latent feature distributions.
Quantitative Results. We present our results in Table 4. Our proposed NexusFlow significantly outperforms all partially supervised baselines. Although still left behind with the Full-supervision, it has a strong advantage in computational cost, especially when compared with MTPSL [31] in terms of training time and GPU memory (shown in Table 5).